# Seeing the Bigger Picture: 3D Latent Mapping for Mobile Manipulation Policy Learning

### v1

@@ Abstract
footnotetext: We gratefully acknowledge support from NSF CCF-2402689 (ExpandAI), ONR N00014-23-1-2353, and the Technology Innovation Program (20018112, Development of autonomous manipulation and gripping technology using imitation learning based on visual and tactile sensing) funded by the Ministry of Trade, Industry & Energy (MOTIE), Korea.footnotetext: Sunghwan Kim, Zhirui Dai, Dwait Bhatt, and Nikolay Atanasov are with the Department of Electrical and Computer Engineering, University of California San Diego, La Jolla, CA 92093, USA, e-mail: {suk063, zhdai, dhbhatt, natanasov}@ucsd.edu.footnotetext: Woojeh Chung, Arth Shukla, and Hao Su are with the Department of Computer Science and Engineering, University of California San Diego, La Jolla, CA 92093, USA, e-mail: {w5chung, arshukla, haosu}@ucsd.edu.footnotetext: Yulun Tian is with the Robotics Department, University of Michigan, Ann Arbor, MI 48109, USA, e-mail: yulunt@umich.edu.
@@ I Introduction
A key idea in this paper is to condition the robot policy on global features obtained from a persistent 3D map (see LABEL:fig:teaser), built incrementally from streaming observations. Persistent maps have long benefited robot navigation [14, 15], yet their potential to enhance manipulation remains under-explored. Such maps offer two key advantages for mobile manipulation. First, they act as spatial memory, offering global visibility of object locations and task goals while mitigating occlusions from the current field of view. Second, by accumulating information over time, they provide long-term context that enables policies to reason beyond short observation windows. LABEL:fig:teaserb shows a latent map containing spatially grounded language features, generated by our method. LABEL:fig:teaserc illustrates that a policy conditioned on the map may attend to the entire scene during task execution and leverage spatially and temporally extended context.
Contributions. We present Seeing the Bigger Picture (SBP), an end-to-end policy learning method operating directly on an incrementally constructed 3D latent feature map. We propose a modular design that comprises (i) a scene-specific latent feature grid that aggregates and compresses multiview visual observations, and (ii) a pre-trained, scene-agnostic decoder that reconstructs target embeddings (e.g., CLIP [16]) from the latent features and generalizes to diverse scenes. During task execution, our approach enables online optimization of latent features by using the pre-trained decoder. To harness the 3D latent map in mobile manipulation tasks, we propose a 3D feature aggregator that summarizes the spatially distributed features into a compact global map token that provides global context to support robot policy learning. The global map token can be integrated as an additional input into a behavior cloning or reinforcement learning policy, and we show that the resulting policy achieves improvements in long-horizon reasoning and task execution.
Our contributions are summarized as follows.
@@ II Related Work
A growing body of work distills vision model features, e.g., CLIP [16] and DINOv2 [17], into neural scene representations for grounding semantics in 3D. Early work such as LERF [18] embeds vision-language features into implicit neural fields by enforcing multiview embedding consistency, enabling open-vocabulary queries. To achieve fine-grained segmentation, subsequent works adopt explicit 3D representations (e.g., Gaussian splatting [19]) and often leverage SAM [20] to obtain precise object boundaries [21, 22]. Recent work extends these representations to dynamic environments through online updates [23] and 4D scene representations for moving objects [24, 25].
Several works structure the 3D features to support downstream manipulation tasks, such as 6-DoF grasping. An early method, F3RM [12], distills CLIP features into a hierarchical 3D feature grid, enabling few-shot grasping by optimizing grasp poses based on language queries. LERF-TOGO [11] addresses non-uniform object activations in LERF [18] using DINO-based object masking, leading to improved part-aware grasp selection. GeFF [26] eliminates the need for per-scene optimization by introducing a generalizable NeRF [27] encoder that predicts language-aligned features. OK-Robot [28] presents a modular pipeline for language-conditioned household manipulation, while DynaMem [29] utilizes a dynamic 3D voxel map with semantic vectors for object localization and navigation. More recently, UAD [30] distills visual affordance maps from VLMs into a task-conditioned predictor that generalizes in the wild.
Despite their strong 3D grounding capabilities, these methods primarily focus on localizing objects or predicting affordances from language queries rather than leveraging the representations for end-to-end policy learning. To bridge this gap, we introduce a 3D latent mapping approach explicitly designed to facilitate policy training for long-horizon tasks.
@@ II-B Manipulation Policy Learning with 3D Representations
One line of work lifts features from 2D large vision models into 3D structures. Early work PerAct [31] encodes RGB-D observations into voxel grids and uses a transformer to predict the next end-effector pose [31]. In a related advance, Act3D [5] lifts multi-scale CLIP features into a 3D feature cloud and scores candidate points via cross-attention to regress manipulator poses. This approach has been extended to generate full action trajectories by conditioning diffusion models on the featurized 3D scene [6]. Alternatively, GNFactor [7] trains a neural feature field that jointly reconstructs 2D foundation-model features and predicts robot actions in 3D.
Another line of work bypasses 2D feature lifting and learns policies directly from geometric observations like point clouds. DP3 encodes raw point clouds into compact embeddings that condition a diffusion policy for manipulation  [3, 9]. To improve precision, some approaches render the scene and apply multi-stage transformers that progressively zoom in on regions of interest [32]. To reintroduce rich semantics, recent work fuses point clouds with DINO features, forming semantic fields that enable part-level generalization across object instances [8]. Leveraging 3D geometric priors, EquiBot combines $\mathrm{SIM}(3)$-equivariant architectures with a point cloud diffusion policy, achieving invariance to scale, rotation, and translation [10]. ActionFlow uses a transformer that attends to $\mathrm{SE}(3)$-invariant point relations and leverages flow matching to learn symmetry-preserving trajectories [33].
Despite recent progress, most existing approaches treat 3D scene representations as instantaneous observations, reconstructing them from scratch at each time step. We address these limitations by learning end-to-end manipulation policies conditioned on a persistent 3D latent map, enabling global and long-horizon reasoning.
@@ III-B Map-Conditioned Policy Learning
We consider a manipulator that executes language-specified instructions. The task specification is embedded in a vector representation $e_{\ell}$ described below. At each time step $\tau$, the robot receives its proprioceptive state $s_{\tau}$ (e.g., joint angles) and observations $o_{\tau}$ (e.g., RGB images). A key feature of our formulation is the 3D latent map $\mathcal{M}_{\tau}$, which is built incrementally from the observations and serves as persistent spatial memory of the environment (see LABEL:fig:teaser). A policy $\pi_{\phi}$ with parameters $\phi$ maps the latent map $\mathcal{M}_{\tau}$, the current state $s_{\tau}$, the observation $o_{\tau}$, and the task embedding $e_{\ell}$ to an action $a_{\tau}$ (e.g., joint velocities).
@@ IV-B Latent Map Optimization
During task execution, the pre-trained decoder can be kept frozen. Adaptation to new environments is accelerated by focusing optimization on the grid parameters $\psi$. The parameters $\psi$ and, optionally, $\theta$ are optimized by minimizing the loss $\mathcal{L}$ in (1). We align the predicted feature $\hat{y}(x)$ with the reference $y$ using the cosine similarity loss, which empirically outperforms alternatives such as the $L_{2}$ loss. Given a dataset $\mathcal{D}$ of coordinate-feature pairs $(x,y)$, the objective is:
@@ V-A Global Map Token
First, we extract features from the map at the vertices of its finest grid level. Let $\mathcal{Z}_{L}=\{z_{L,i}\}_{i=1}^{M_{L}}$ be the coordinates of these vertices. We compute a decoded feature at each vertex using the multiresolution feature $F_{\psi}(z_{L,i})$ from (5) and the decoder $D_{\theta}$:
$\hat{y}(z_{L,i})=D_{\theta}\bigl(F_{\psi}(z_{L,i})\bigr).$ (8)
We then form the set $\mathcal{S}_{L}$ of coordinate–feature pairs, using only vertices marked as occupied:
@@
where $v_{i}\in\{0,1\}$ is an occupancy indicator for vertex $z_{L,i}$. The features are processed by a 3D feature aggregator, $E_{\mathrm{3D}}$, whose output is max-pooled to produce the global map token:
$e_{m}=\operatorname{Max-Pool}\Bigl(E_{\mathrm{3D}}(\mathcal{S}_{L})\Bigr).$ (10)
@@
Figure 4: Global map token. Latent features from $\mathcal{F}$ are decoded to $\mathcal{Y}$ via the decoder $D_{\theta}$ at the finest grid vertices. The 3D feature aggregator processes the coordinate-feature pairs, and its output is max-pooled to produce the global map token $e_{m}$.
The $E_{\mathrm{3D}}$ architecture matches the environment’s scale. For large room-scale environments, we use a Point Transformer [40], whose hierarchical attention mechanism is well-suited for capturing long-range spatial relationships. For smaller tabletop scenes, we utilize a lightweight PointNet [41], whose simpler structure is sufficient for compact environments and its computational efficiency is advantageous for sample-intensive RL. We encode coordinates with sinusoidal positional encodings [27], concatenate them with features, and input them to PointNet.
@@ V-B Map-Conditioned Policy Network
We treat the map token $e_{m}$ in (10) as an additional state input to the policy networks. We first outline a generic integration scheme and, then, specify its instantiation for BC and RL. The inputs to the policy network consist of image features $E_{I}(o_{\tau})$, proprioceptive state $s_{\tau}$, task embedding $e_{\ell}$, and the global map token $e_{m}$. Concatenating these inputs yields a joint embedding $h_{\tau}$:
@@ VI Evaluation
Across all experiments, performance is evaluated using success rate (SR) and episode reward (ER). The composition of ER varies by task but is composed of object reach shaping, a grasp bonus, and a success bonus. We evaluate performance on both training settings (in-distribution, ID) and novel settings (out-of-distribution, OOD). All simulations are carried out in the ManiSkill simulator [43].
@@ VI-A Mobile Manipulation Experiment
TABLE II: Performance of RL policies on the sequential pick-and-place task. We report success rate (SR↑) and episode reward (ER↑), averaged over three runs. Best results are highlighted in bold.
@@ VI-B Sequential Manipulation Experiment
Setup. We designed a two-stage sequential pick-and-place task to evaluate whether the latent map improves performance on long-horizon manipulation. The task requires picking objects from a cluttered tabletop and placing them in a basket in a prescribed order (Fig. 7). For each episode, we sample an environment from a predefined set of scenes that differ in the arrangements and poses of clutter and target objects. The target set is fixed across episodes, and the execution order of targets is randomized per episode. To mimic limited visibility in mobile manipulation, the robot relies solely on an egocentric camera (i.e., no global view of the scene). The scenes contain occlusions, making it challenging to localize the target objects. The experiments are conducted on a uFactory xArm6 robot. We evaluate on training (ID) scenes and novel (OOD) scenes with unseen object arrangements.
@@ VII Conclusion and Future Work
Several avenues for future work are possible. First, while our latent map provides global, long-term context, the policy still relies on local image features. Future work could reduce this dependency by developing dynamic scene representations to capture robot arm and other object motion. Second, the map-conditioned RL policy is trained using an on-policy method, which can be sample-inefficient, requiring pre-training from an image-based policy. Integrating our method with off-policy or model-based RL could mitigate these issues. Finally, evaluating our approach on larger-scale, longer-horizon mobile manipulation tasks would further validate its effectiveness in diverse real-world settings.

### v3

@@ Abstract
footnotetext: We gratefully acknowledge support from NSF CCF-2402689 (ExpandAI), ONR N00014-23-1-2353, and the Technology Innovation Program (20018112, Development of autonomous manipulation and gripping technology using imitation learning based on visual and tactile sensing) funded by the Ministry of Trade, Industry & Energy (MOTIE), Korea.footnotetext: S. Kim, Z. Dai, D. Bhatt, and N. Atanasov are with the Department of Electrical and Computer Engineering, University of California San Diego, La Jolla, CA 92093, USA, e-mail: {suk063, zhdai, dhbhatt, natanasov}@ucsd.edu.footnotetext: W. Chung, A. Shukla, and H. Su are with the Department of Computer Science and Engineering, University of California San Diego, La Jolla, CA 92093, USA, e-mail: {w5chung, arshukla, haosu}@ucsd.edu.footnotetext: Y. Tian is with the Robotics Department, University of Michigan, Ann Arbor, MI 48109, USA, e-mail: yulunt@umich.edu.
@@ I Introduction
A key idea in this paper is to condition the robot policy on global features obtained from a persistent 3D map (see LABEL:fig:teaser), built incrementally from streaming observations. Persistent maps have long benefited robot navigation [5, 2], yet their potential to enhance manipulation remains under-explored. Such maps offer two key advantages for mobile manipulation. First, they act as spatial memory, offering global visibility of object locations and task goals while mitigating occlusions from the current field of view. Second, by accumulating information over time, they provide long-term context that enables policies to reason beyond short observation windows. LABEL:fig:teaserb shows a latent map containing spatially grounded language features, generated by our method. LABEL:fig:teaserc illustrates that a policy conditioned on the map may attend to the entire scene, leveraging spatially and temporally extended context.
Contributions. We present Seeing the Bigger Picture (SBP), an end-to-end policy learning method operating directly on an incrementally constructed 3D latent feature map. We propose a modular design that comprises (i) a scene-specific latent feature grid that aggregates and compresses multiview visual observations, and (ii) a pre-trained, scene-agnostic decoder that reconstructs target embeddings (e.g., CLIP [27]) from the latent features and generalizes to diverse scenes. During task execution, our approach enables online optimization of latent features by using the pre-trained decoder. To harness the 3D latent map in mobile manipulation tasks, we propose a 3D feature aggregator that summarizes the spatially distributed features into a compact global map token that provides global context to support robot policy learning. The global map token can be integrated into a behavior cloning or reinforcement learning policy, and we show that the resulting policy improves long-horizon reasoning and task execution. Our contributions are summarized as follows.
@@ II Related Work
A growing body of work distills vision model features, e.g., CLIP [27] and DINOv2 [21], into neural scene representations for grounding semantics in 3D. Early work such as LERF [14] embeds vision-language features into implicit neural fields by enforcing multiview consistency, enabling open-vocabulary queries. To achieve fine-grained segmentation, subsequent works adopt explicit 3D representations (e.g., Gaussian splatting [13]) and leverage SAM [15] to obtain precise object boundaries [24, 26]. Recent work extends these representations to dynamic environments via online updates [11] and 4D scene representations for moving objects [39, 16].
Several works structure the 3D features to support downstream manipulation tasks, such as 6-DoF grasping. F3RM [30] distills CLIP features into a hierarchical 3D feature grid, enabling few-shot grasping by optimizing grasp poses based on language queries. LERF-TOGO [28] addresses non-uniform activations in LERF [14] via DINO-based masking, improving part-aware grasp selection. GeFF [25] eliminates the need for per-scene optimization by introducing a generalizable NeRF [19] encoder that predicts language-aligned features. OK-Robot [18] presents a modular pipeline for language-conditioned household manipulation, while DynaMem [17] utilizes a dynamic 3D voxel map with semantic vectors for object localization and navigation. More recently, UAD [35] distills visual affordance maps from VLMs into a task-conditioned predictor that generalizes in the wild.
Despite their strong 3D grounding capabilities, these methods focus on localizing objects or predicting affordances from language queries rather than leveraging the representations for end-to-end policy learning. To bridge this gap, we introduce a 3D latent mapping approach designed to facilitate policy training for long-horizon tasks.
@@ II-B Manipulation Policy Learning with 3D Representations
One line of work lifts features from 2D large vision models into 3D structures. PerAct [31] encodes RGB-D observations into voxel grids and predicts end-effector poses via a transformer. Act3D [8] lifts multi-scale CLIP features into a 3D feature cloud and scores candidate points via cross-attention to regress manipulator poses. This has been extended to generate action trajectories by conditioning diffusion models on the featurized 3D scene [12]. GNFactor [41] trains a neural feature field that jointly reconstructs 2D foundation-model features and predicts robot actions in 3D.
Another line of work bypasses 2D feature lifting and learns policies directly from geometric observations like point clouds. DP3 encodes raw point clouds into compact embeddings that condition a diffusion policy for manipulation [3, 42]. Some approaches render the scene and apply multi-stage transformers that zoom in on regions of interest [9]. Recent work fuses point clouds with DINO features, forming semantic fields that enable part-level generalization across object instances [37]. EquiBot [40] combines $\mathrm{SIM}(3)$-equivariant architectures with a point cloud diffusion policy for scale, rotation, and translation invariance, while ActionFlow [7] leverages $\mathrm{SE}(3)$-invariant point relations with flow matching for symmetry-preserving trajectories.
Despite recent progress, most approaches treat 3D representations as instantaneous observations, reconstructing them at each step. We address this by learning end-to-end policies conditioned on a persistent 3D latent map, enabling global and long-horizon reasoning.
@@ III-B Map-Conditioned Policy Learning
We consider a manipulator that executes language-specified instructions. The task specification is embedded in a vector $e_{\ell}$ described below. At each time step $\tau$, the robot receives its proprioceptive state $s_{\tau}$ (e.g., joint angles) and observations $o_{\tau}$ (e.g., RGB images). A key feature of our formulation is the 3D latent map $\mathcal{M}_{\tau}$, which is built incrementally from the observations and serves as persistent spatial memory of the environment (see LABEL:fig:teaser). A policy $\pi_{\phi}$ with parameters $\phi$ maps $\mathcal{M}_{\tau}$, state $s_{\tau}$, the observation $o_{\tau}$, and the task embedding $e_{\ell}$ to an action $a_{\tau}$ (e.g., joint velocities).
@@ IV-B Latent Map Optimization
During task execution, the pre-trained decoder can be kept frozen. Adaptation to new environments is accelerated by focusing optimization on the grid parameters $\psi$. The parameters $\psi$ and, optionally, $\theta$ are optimized by minimizing the loss $\mathcal{L}$ in (1). We align the predicted feature $\hat{y}(x)$ with the reference $y$ using the cosine distance loss, which empirically outperforms alternatives such as the $L_{2}$ loss. Given a dataset $\mathcal{D}$ of coordinate-feature pairs $(x,y)$, the objective is:
@@ V-A Global Map Token
First, we extract features from the map at the vertices of its finest grid level. Let $\mathcal{Z}_{L}=\{z_{L,i}\}_{i=1}^{M_{L}}$ be the coordinates of these vertices. We compute a decoded feature at each vertex using the multiresolution feature $F_{\psi}(z_{L,i})$ from (5) and the decoder $D_{\theta}$: $\hat{y}(z_{L,i})=D_{\theta}\bigl(F_{\psi}(z_{L,i})\bigr)$. We then form the set $\mathcal{S}_{L}$ of coordinate–feature pairs at selected vertices:
@@
where $v_{i}\in\{0,1\}$ is a binary mask selecting task-relevant vertices (e.g., excluding unoccupied or background regions). These are fed into a 3D feature aggregator, $E_{\mathrm{3D}}$, and its outputs are attention-pooled to produce a global map token:
$e_{m}=\operatorname{Attn-Pool}\Bigl(E_{\mathrm{3D}}(\mathcal{S}_{L})\Bigr).$ (9)
@@
Figure 4: Global map token. Latent features from $\mathcal{F}$ are decoded to $\mathcal{Y}$ via the decoder $D_{\theta}$ at the finest grid vertices. The 3D feature aggregator processes the coordinate-feature pairs, and its outputs are attention-pooled to produce the global map token $e_{m}$.
The $E_{\mathrm{3D}}$ architecture matches the environment’s scale. For large room-scale environments, we use a Point Transformer [43], whose hierarchical attention captures long-range spatial relationships. For smaller tabletop scenes, we utilize a lightweight PointNet [23], whose simpler structure is sufficient for compact environments and its computational efficiency is advantageous for sample-intensive RL. We encode coordinates with sinusoidal positional encodings [19], concatenate them with features, and input them to PointNet.
@@ V-B Map-Conditioned Policy Network
We treat the map token $e_{m}$ in (9) as an additional state input to the policy networks. We first outline a generic integration scheme and, then, specify its instantiation for BC and RL. The policy inputs are image features $E_{I}(o_{\tau})$, proprioceptive state $s_{\tau}$, task embedding $e_{\ell}$, and the global map token $e_{m}$. Concatenating these inputs yields a joint embedding $h_{\tau}$:
@@ VI Evaluation
Across all experiments, performance is evaluated using success rate (SR) and episode reward (ER). ER comprises object reach shaping, a grasp bonus, and a success bonus. We report both in-distribution (ID) and out-of-distribution (OOD) results. All experiments run in the ManiSkill simulator [36].
@@ VI-A Mobile Manipulation Experiment
TABLE II: Performance of RL policies on the sequential pick-and-place task. We evaluate on 100 training (ID) and 20 novel (OOD) scenes. We report success rate (SR↑) and episode reward (ER↑), averaged over three runs. Best results are highlighted in bold.
@@ VI-B Sequential Manipulation Experiment
Setup. We design a two-stage sequential pick-and-place task to evaluate whether the latent map improves long-horizon manipulation. The task requires picking objects from a cluttered tabletop and placing them in a basket in a prescribed order (Fig. 7). For each episode, we sample an environment from a predefined set of scenes that differ in the arrangements and poses of clutter and target objects. The target set is fixed across episodes, and the execution order of targets is randomized per episode. To mimic limited visibility in mobile manipulation, the robot relies solely on an egocentric camera (i.e., no global view of the scene). The scenes contain occlusions, making it challenging to localize the target objects. Experiments are conducted on a uFactory xArm6 robot. We evaluate on 100 training (ID) scenes and 20 novel (OOD) scenes with unseen object arrangements.
@@ VII Conclusion
Several avenues for future work are possible. First, while our latent feature map provides global context, the policy still relies on local image features. Future work could reduce this dependency by developing dynamic scene representations to capture the motion of the robot and other objects. Second, the map-conditioned RL policy is trained using an on-policy method, which can be sample-inefficient, requiring pre-training from an image-based policy. Integrating our method with off-policy or model-based RL could mitigate these problems. Third, our map-conditioned policy uses Point Transformer [43] for behavior cloning but a lighter PointNet [23] for RL, as the former is too expensive for online training. Developing a unified and efficient 3D aggregation model would enable online policy training. Finally, evaluating our approach on larger-scale, longer-horizon mobile manipulation tasks is needed to further validate its effectiveness. Extending to more complex dexterous manipulation scenarios could also demonstrate its broader applicability.